a man walks from his home at a speed of 5km/h and returns at a speed of 3km/h what is the distance he walked if the whole trip takes 4 hours
Answer by ankor@dixie-net.com(22740)

a man walks from his home at a speed of 5km/h and returns at a speed of 3km/h what is the distance he walked if the whole trip takes 4 hours
:
Find the time each way first
t = 5 mph time
then
(4-t) = 3 mph time
:
Write a time equation; time = speed * time
:
Go dist = return dist
5t = 3(4-t)
5t = 12 - 3t
5t + 3t = 12
t = 12/8
t = 1.5 hrs
:
Find the distance
5(1.5) = 7.5 mi
check the dist
3(4-1.5) = 7.5 mi